# Enhancing NMR Shielding Predictions of Atoms-in-Molecules Machine Learning Models with Neighborhood-Informed Representations

### v1

@@ III.1.1 Atomic Coulomb matrix (aCM)
The CM is an $N_{\rm atm}\times N_{\rm atm}$ symmetric matrix, where $N_{\rm atm}$ is the number of atoms in a molecule. Its elements are defined as follows: off-diagonal terms are given by Coulombic interactions, while diagonal terms are polynomial fits of the nuclear charge (i.e. atomic number) $Z$ to the free atom’s total energy Rupp et al. (2012):
@@
where $Z_{I}$ is the atomic number and $R_{IJ}$ is the interatomic distance between nuclei $I$ and $J$. The CM is invariant to translation, rotation, and reflection, but not to the permutation of atomic indices. To address this, row-sorted, column-sorted, eigenspectrum-based, and randomly permuted CMs have been introduced Hansen et al. (2013). Barker et al. further localized CM representations for use in Gaussian approximation potentials (GAP), resulting in localized-CM, decaying-CM, and reduced-CM variants Barker et al. (2017).
To represent atomic environments, the aCM can be constructed. Tripathy et al. employed aCM descriptors to predict $1s$ core-electron binding energies of small organic molecules Tripathy et al. (2024a). In aCM, the target atom is placed in the first index, and the remaining atoms are ordered by increasing distance from it.
@@ III.1.3 Continuous atomic Coulomb matrix (aCM-RBF)
RBFs such as Gaussians are widely used in molecular descriptors, for example, in SLATM Huang and von Lilienfeld (2020) and in Fourier-series expansions of atomic radial distribution functions Von Lilienfeld et al. (2015). The resulting continuous representation, denoted aCM-RBF, is given by
@@
III.1.4 Continuous bag-of-bonds (aBoB-RBF)
@@
III.1.5 Atomic Spectrum of London Axilrod–Teller–Muto (aSLATM)
The SLATM descriptor, and its atomic version aSLATM, encode molecular environments using one-, two-, and three-body terms Huang and von Lilienfeld (2020). For an atom, $I$, the one-body contribution is the atomic number:
@@ III.1.5 Atomic Spectrum of London Axilrod–Teller–Muto (aSLATM)
For the QM9 dataset, $A,B,C\in\{\mathrm{H},\mathrm{C},\mathrm{N},\mathrm{O},\mathrm{F}\}$, giving 5 unique atom types, 15 unique $(A,B)$ pairs, and 35 unique $(A,B,C)$ triplets. A common implementation of SLATM and aSLATM is available in the QML package Christensen et al. (2017).
@@
${\bf d}^{(A,B,C)}(t)=\frac{1}{3}Z_{I}\sum_{J\neq K\neq I}Z_{J}Z_{K}\cdot g_{IJK}(t)\cdot h(t,{\bf R}_{IJ},{\bf R}_{IK}),$ (17)
@@
III.1.6 Neighbor atom information encoded descriptors: aCM($n$), aBoB($n$), aCM-RBF($n$), aBoB-RBF($n$), and aSLATM($n$)
@@
where $d_{k}$ is the distance between the query atom and its $k^{\text{th}}$ neighbor, and $\lambda$, $p$, and $r_{\rm cut}$ are parameters controlling the decay of the respective damping functions. The performance of different damping schemes was benchmarked on the 5k subset of QM9, and for the cosine damping function with $r_{\rm cut}=2.0$ Å, low errors were obtained consistently across all various descriptors as shown in Table S6. Hence, this cos@[$\cos@[2.0]$ scheme is adopted as the default damping function for all descriptors in this study.
Figure 3: Encoding of neighborhood information in atomic descriptors illustrated for the carbonyl carbon (C2) of acetaldehyde. The query atom’s descriptor is ${\bf d}(0)$, which is concatenated with the descriptors of the neighboring atoms after damping with the cos@[$\cos@[2.0]$ function.
@@ III.2 Kernel-Ridge Regression
Kernel-ridge regression is a regularized least squares regression method Kung (2014), and one of the most accurate ML frameworks for modeling chemical properties Rupp et al. (2012). It has been successfully applied to global molecular properties such as atomization energies Rupp et al. (2012); Hansen et al. (2015), as well as local properties including NMR shielding Gupta, Chakraborty, and Ramakrishnan (2021) and core electron binding energies Tripathy et al. (2024a); Porcelli et al. (2025).
@@
where $\sigma$ denotes the kernel width hyperparameter. In earlier studies on molecular global properties, the Laplacian kernel was found to perform well with discrete descriptors Hansen et al. (2015). More recently, Shiota et al. reported that the relative performance of Laplacian and Gaussian kernels in KRR depends on dataset sampling and training set size Shiota, Ishihara, and Mizukami (2024). To assess this in the present context, we evaluated both kernels across all descriptor types for predicting 13C chemical shifts (see Table S8 in the SI). In every case, the Laplacian kernel yielded lower prediction errors, which we use in all KKR models reported in this study.
In the KRR framework, two key hyperparameters that control the accuracy of the model are the regularization strength $\lambda$ (Eq. 25) and the kernel width $\sigma$ (Eq. 30). The regularization parameter $\lambda$ ensures the non-singularity of the kernel matrix and penalizes regression coefficients of large magnitudes to mitigate overfitting. Previous studies recommend small values of $\lambda$ to mitigate linear dependency issues for large training set sizes Ramakrishnan and von Lilienfeld (2015); Tripathy et al. (2024a); Gupta, Chakraborty, and Ramakrishnan (2021). In this study, we set $\lambda=10^{-5}$. The kernel width $\sigma$ controls the spread of similarity across the training set. A heuristic strategy for optimizing $\sigma$ sets the minimum kernel value to $K_{ij}^{\rm min}=0.5$ over the kernel range $K_{ij}\in[0,1]$, yielding Ramakrishnan and von Lilienfeld (2015):
@@
In the QM9NMR study Gupta, Chakraborty, and Ramakrishnan (2021), improved accuracy was obtained by replacing $D_{ij}^{\rm max}$ with the median descriptor difference $D_{ij}^{\rm median}$. This approach was later generalized for core-electron binding energy modeling Tripathy et al. (2024a) by scanning $K_{ij}^{\rm min}$ over a continuous range $\tau\in(0,1)$, leading to the definition of a $\tau$-plot” for kernel width optimization:
@@
In this study, we calculated $D_{ij}^{\rm median}$ using the training entries. This quantity can also be determined using a fixed, reasonably large training set (say, of size 1k) and applied to larger sets. As shown in Figure S3 of the SI, out-of-sample errors are minimized for $\tau\approx 0.3-0.5$ for more accurate descriptors such as aBoB-RBF(4) and aSLATM(0). Accordingly, we adopt $\tau=0.5$ in this study, as this value is consistent with the previous KRR-modeling of NMR shielding Gupta, Chakraborty, and Ramakrishnan (2021).
@@ IV Results and Discussions
Figure 4: a) Learning curves on the log-log scale showing mean absolute errors (MAEs in ppm) for predicting 13C-NMR $\sigma_{\rm iso}$ values with increasing training set sizes. MAEs are calculated for 50k out-of-sample atoms in the QM9NMR dataset. All descriptors have information of the four nearest neighbor atoms (i.e., ${\bf d}(4)$). b) MAEs for a train:test split of 100k:50k for various descriptors with information of $n$ nearest atoms encoded. Previously reported values of 1.88 (FCHL, Ref. Gupta, Chakraborty, and Ramakrishnan, 2021) and 1.87 (MACE-OFF23-small, Ref. Shiota, Ishihara, and Mizukami, 2024) are shown as blue and red horizontal lines. Values for the descriptors studied in this work are shown for various $n$: aCM (maroon), aBoB (violet), aCM-RBF (orange), aBoB-RBF (skyblue), and aSLATM (green).
@@ IV.1 Prediction accuracy on QM9 molecules
To assess how well the ML models capture the chemical environment with increasing data, we plot $\loglog\log$ learning curves of mean absolute error (MAE) versus training set size in Figure 4a. The monotonically decreasing MAEs for aCM(4), aCM-RBF(4), aBoB(4), and aBoB-RBF(4) confirm that local atomic representations effectively encode the chemical environment and improve predictions of $\delta(^{13}{\rm C})$ as training data grows. Notably, the transition from the discrete aCM(4) to the continuous aCM-RBF(4) yields a pronounced reduction in error, whereas for aBoB(4) versus aBoB-RBF(4), the improvement is smaller but still noticeable, indicating that continuous representations capture shielding information more efficiently.
@@
FCHL 1.88 4.2 3.7 Ref. Gupta, Chakraborty, and Ramakrishnan,2021
MACE-OFF23-small 1.87 3.85 2.83 Ref. Shiota, Ishihara, and Mizukami,2024
@@
Overall, aBoB-RBF(4), with an MAE of 1.69 ppm, surpasses the previously reported benchmarks of 1.88 ppm for FCHL Gupta, Chakraborty, and Ramakrishnan (2021) and 1.87 ppm for the MACE-OFF23-small descriptor Shiota, Ishihara, and Mizukami (2024).
Figure 6: Example of predicting 13C-NMR chemical shifts using mlqm9nmr module mlq (2025). For more detail, see mlqm9nmr module in SI.
Figure 7: DFT calculated (NMR: mPW1PW91/6-311+G(2d,p), geometry: B3LYP/6-31G(2df,p)) and ML-predicted 13C-NMR spectra for randomly chosen molecules, each from Drug12, Drug40, pyrimidineone, and GDBm validation set. ML predictions were made with mlqm9nmr mlq (2025) using aBob-RBF(4) descriptor generated with PM7 geometries. See Figure 6 for an example query and SI for further details.

### v3

@@ III.1.1 Atomic Coulomb matrix (aCM)
The CM is an $N_{\rm atm}\times N_{\rm atm}$ symmetric matrix, where $N_{\rm atm}$ is the number of atoms in a molecule. Its elements are defined as follows: off-diagonal terms are given by Coulombic interactions, while diagonal terms are polynomial fits of the nuclear charge (i.e. atomic number) $Z$ to the free atom’s total energy rupp2012fast:
@@
where $Z_{I}$ is the atomic number and $R_{IJ}$ is the interatomic distance between nuclei $I$ and $J$. The CM is invariant to translation, rotation, and reflection, but not to the permutation of atomic indices. To address this, row-sorted, column-sorted, eigenspectrum-based, and randomly permuted CMs have been introduced hansen2013assessment. Barker et al. further localized CM representations for use in Gaussian approximation potentials (GAP), resulting in localized-CM, decaying-CM, and reduced-CM variants Barker2017.
To represent atomic environments, the aCM can be constructed. Tripathy et al. employed aCM descriptors to predict $1s$ core-electron binding energies of small organic molecules tripathy2024chemical. In aCM, the target atom is placed in the first index, and the remaining atoms are ordered by increasing distance from it.
@@ III.1.3 Continuous atomic Coulomb matrix (aCM-RBF)
RBFs such as Gaussians are widely used in molecular descriptors, for example, in SLATM huang2020quantum and in Fourier-series expansions of atomic radial distribution functions von2015fourier. The resulting continuous representation, denoted aCM-RBF, is given by
@@
III.1.4 Continuous atomic bag-of-bonds (aBoB-RBF)
@@
III.1.5 Physical motivation for multiplying the Coulomb matrix with a scaling function
The scaling factor $s(R_{IJ})=1/R_{IJ}^{\beta}$ (see Eq. 3) is not merely a numerical device for suppressing distant atomic contributions. It has a direct physical motivation in the long-range behavior of NMR shielding, $\sigma(R)$. The NMR shielding tensor is a second-order magnetic response property that couples the external magnetic field to induced electronic currents. Analyses of atom–atom interactions by Barszczewicz et al. barszczewicz1996long showed that the deviation of the shielding of an interacting atom, from the free-atom value, decays asymptotically as $1/R^{n}$, with $n=3,4,$ and $6$ for the anisotropy of the shielding tensor (dipolar contribution), atom–ion interactions (induction-dominated), and the isotropic shielding of two neutral atoms (dispersion-dominated), respectively. These inverse-power laws reflect the fact that shielding is controlled not by static electrostatics but by induced magnetic fields and current-density response, which decay much more rapidly with distance. Thus, a descriptor whose radial decay lies within the physically relevant range $1/R^{3}$$1/R^{6}$ is far more consistent with magnetic response theory than an undamped Coulomb matrix suitable to map to molecular total energy.
Multiplying the Coulomb interaction $Z_{I}Z_{J}/R_{IJ}$ by an additional factor of $1/R_{IJ}^{3}$ yields descriptor elements proportional to placing the effective decay rate between the dominant long-range regimes. Importantly, for typical organic molecules such as those in QM9, the electronic environments around heteroatoms and substituted carbons resemble quasi-ion or quasi-dipole centers rather than spherically symmetric neutral atoms. Local bond polarity, electronegativity differences, and hybridization generate partial charges and permanent or induced dipoles. In such situations, the induction mechanism, which leads to an $1/R^{4}$ scaling of the isotropic shielding, is expected to dominate over the dispersion-driven $1/R^{6}$ decay characteristic of strictly neutral closed-shell atoms. Hence, the $1/R^{4}$ behavior implicit in the scaled Coulomb matrix is not only physically justified but closely aligned with the shielding response expected for realistic molecular environments.
III.1.6 Atomic Spectrum of London Axilrod–Teller–Muto (aSLATM)
The SLATM descriptor, and its atomic version aSLATM, encode molecular environments using one-, two-, and three-body terms huang2020quantum. For an atom, $I$, the one-body contribution is the atomic number:
@@ III.1.6 Atomic Spectrum of London Axilrod–Teller–Muto (aSLATM)
For the QM9 dataset, $A,B,C\in\{\mathrm{H},\mathrm{C},\mathrm{N},\mathrm{O},\mathrm{F}\}$, giving 5 unique atom types, 15 unique $(A,B)$ pairs, and 35 unique $(A,B,C)$ triplets. A common implementation of SLATM and aSLATM is available in the QML package QML.
@@
${\bf d}_{I}^{(A,B,C)}(t)=\frac{1}{3}Z_{I}\sum_{J\neq K\neq I}Z_{J}Z_{K}\cdot g_{IJK}(t)\cdot h(t,{\bf R}_{IJ},{\bf R}_{IK}),$ (17)
@@
III.1.7 Neighbor atom information encoded descriptors: aCM($n$), aBoB($n$), aCM-RBF($n$), aBoB-RBF($n$), and aSLATM($n$)
@@
where $d_{k}$ is the distance between the query atom and its $k^{\text{th}}$ neighbor, and $\lambda$, $p$, and $r_{\rm cut}$ are parameters controlling the decay of the respective damping functions. The performance of different damping schemes was benchmarked on the 5k subset of QM9, and for the cosine damping function with $r_{\rm cut}=2.0$ Å, low errors were obtained consistently across all various descriptors as shown in Table S6. Hence, this $f_{\rm cos}(d_{k};2.0)$ scheme is adopted as the default damping function for all descriptors in this study. The implementation of the damped neighbour descriptors is described below. For example, in the C2 atom in acetaldehyde as shown in Figure 3, with $n=0$, the descriptor is ${\bf d}(0)=\left[{\bf d}_{\rm C2}\right]$. For $n=1$, with inclusion of the damping function, the descriptor becomes ${\bf d}(1)=\left[{\bf d}_{\rm C2},\,{\bf d}_{\rm H4}\cdot f_{\rm cos}(R_{\rm C2-H4};2.0)\right]$. Similarly, for $n=2$, we arrive at ${\bf d}(2)=\left[{\bf d}_{\rm C2},\,{\bf d}_{\rm H4}\cdot f_{\rm cos}(R_{\rm C2-H4};2.0),\,{\bf d}_{\rm O1}\cdot f_{\rm cos}(R_{\rm C2-O1};2.0)\right]$, and so forth. Here, $R_{\rm C2-H4}$ and $R_{\rm C2-O1}$ denote the distance between the query C2-atom and its first and second neighbor atoms, respectively (see Figure 3).
In summary, for both the query atom and its neighbors, the scaling function $s$ reduces the weight of descriptor elements as the interatomic distance increases, as illustrated in Fig. 2. The damping function $f$ further suppresses the contribution of the entire neighbor descriptor vector for the query atom with increasing distance (Fig. 3). Together, $s$ and $f$ ensure that the descriptor of a query atom remains “size-consistent”, meaning that its representation of the local atomic environment does not change when the overall molecular size increases. This property enables consistent training and prediction across molecules of varying sizes.
Figure 3: Encoding of neighborhood information in atomic descriptors illustrated for the carbonyl carbon (C2) of acetaldehyde. The query atom’s descriptor is ${\bf d}(0)$, which is concatenated with the descriptors of the neighboring atoms after damping with the $f_{\rm cos}(d_{k};2.0)$ function.
@@ III.2 Kernel-Ridge Regression
Kernel-ridge regression is a regularized least squares regression method kung2014kernel, and one of the most accurate ML frameworks for modeling chemical properties rupp2012fast. It has been successfully applied to global molecular properties such as atomization energies rupp2012fast; hansen2015machine, as well as local properties including NMR shielding gupta2021revving and core electron binding energies tripathy2024chemical; porcelli2025photoemission.
@@
where $\sigma$ denotes the kernel width hyperparameter. In earlier studies on molecular global properties, the Laplacian kernel was found to perform well with discrete descriptors hansen2015machine. More recently, Shiota et al. reported that the relative performance of Laplacian and Gaussian kernels in KRR depends on dataset sampling and training set size shiota2024universal. To assess this in the present context, we evaluated both kernels across all descriptor types for predicting 13C chemical shifts (see Table S8 in the SI). In every case, the Laplacian kernel yielded lower prediction errors, which we use in all KKR models reported in this study.
In the KRR framework, two key hyperparameters that control the accuracy of the model are the regularization strength $\lambda$ (Eq. 25) and the kernel width $\sigma$ (Eq. 30). The regularization parameter $\lambda$ ensures the non-singularity of the kernel matrix and penalizes regression coefficients of large magnitudes to mitigate overfitting. Previous studies recommend small values of $\lambda$ to mitigate linear dependency issues for large training set sizes ramakrishnan2015many; tripathy2024chemical; gupta2021revving. In this study, we set $\lambda=10^{-5}$. The kernel width $\sigma$ controls the spread of similarity across the training set. A heuristic strategy for optimizing $\sigma$ sets the minimum kernel value to $K_{ij}^{\rm min}=0.5$ over the kernel range $K_{ij}\in[0,1]$, yielding ramakrishnan2015many:
@@
In the QM9NMR study gupta2021revving, improved accuracy was obtained by replacing $D_{ij}^{\rm max}$ with the median descriptor difference $D_{ij}^{\rm median}$. This approach was later generalized for core-electron binding energy modeling tripathy2024chemical by scanning $K_{ij}^{\rm min}$ over a continuous range $\tau\in(0,1)$, leading to the definition of a $\tau$-plot” for kernel width optimization:
@@
In this study, we calculated $D_{ij}^{\rm median}$ using the training entries. This quantity can also be determined using a fixed, reasonably large training set (say, of size 1k) and applied to larger sets. As shown in Figure S3 of the SI, out-of-sample errors are minimized for $\tau\approx 0.3-0.5$ for more accurate descriptors such as aBoB-RBF(4) and aSLATM(0). Accordingly, we adopt $\tau=0.5$ in this study, as this value is consistent with the previous KRR-modeling of NMR shielding gupta2021revving.
@@ IV Results and Discussions
Figure 4: a) Learning curves on the log-log scale showing mean absolute errors (MAEs in ppm) for predicting 13C-NMR $\sigma_{\rm iso}$ values with increasing training set sizes. MAEs are calculated for 50k out-of-sample atoms in the QM9NMR dataset. All descriptors have information of the four nearest neighbor atoms (i.e., ${\bf d}(4)$). b) MAEs for a train:test split of 100k:50k for various descriptors with information of $n$ nearest atoms encoded. Previously reported values of 1.88 (FCHL, Ref. gupta2021revving) and 1.87 (MACE-OFF23-small, Ref. shiota2024universal) are shown as blue and red horizontal lines. Values for the descriptors studied in this work are shown for various $n$: aCM (maroon), aBoB (violet), aCM-RBF (orange), aBoB-RBF (skyblue), and aSLATM (green).
@@ IV.1 Prediction accuracy on QM9 molecules
To assess how well the ML models capture the chemical environment with increasing data, we plot $\loglog\log$ learning curves of mean absolute error (MAE) versus training set size in Figure 4a. The monotonically decreasing MAEs for aCM(4), aCM-RBF(4), aBoB(4), and aBoB-RBF(4) confirm that local atomic representations effectively encode the chemical environment and improve predictions of $\delta(^{13}{\rm C})$ as training data grows. Notably, the transition from the discrete aCM(4) to the continuous aCM-RBF(4) yields a pronounced reduction in error, whereas for aBoB(4) versus aBoB-RBF(4), the improvement is smaller but still noticeable, indicating that continuous representations capture shielding information more efficiently. To further quantify learning efficiency, we fit the learning curves to the power-law form $\log(\mathrm{Error})=a+b,\log(N)$. The fitted slopes for aCM(4), aBoB(4), aCM-RBF(4), and aBoB-RBF(4) are 0.281-0.281, 0.216-0.216, 0.236-0.236, and 0.260-0.260, respectively, indicating that all four descriptors exhibit comparable learning behavior. In practice, the slope of a learning curve is strongly problem dependent and reflects the underlying complexity of the target function. For molecular machine-learning tasks such as NMR shielding, where the property depends on detailed, high-dimensional local environments, learning curves typically decay more slowly. The slopes observed here (approximately 0.25-0.25, corresponding to $\mathrm{Error}\propto 1/N^{0.25}$) are therefore consistent with expectations for this type of problem, rather than faster decays such as $\mathrm{Error}\propto 1/N$ or $\mathrm{Error}\propto 1/\sqrt{N}$. The strictly monotonic decrease in MAE confirms stable and systematic learning as the dataset grows across all descriptors.
@@
FCHL 1.88 4.2 3.7 Ref. gupta2021revving
MACE-OFF23-small 1.87 3.85 2.83 Ref. shiota2024universal
@@
We also modeled 13C chemical shifts from the QM9NMR dataset using the $\Delta$-ML approach ramakrishnan2015big. When using HF/STO-3G baseline computed on PM7 geometries and the aBoB-RBF(4) descriptor, the MAEs on a 50k hold-out test set decrease to 1.66, 1.46, 1.29, and 1.17 ppm for training set sizes of 16k, 32k, 64k, and 100k, respectively. Using mPW1PW91/STO-3G baseline calculations on the same PM7 geometries yields very similar performance, with MAEs of 1.66, 1.46, 1.28, and 1.16 ppm for the corresponding training sizes. For comparison, Gupta et al. reported an MAE of 1.36 ppm for gas-phase 13C NMR chemical shielding using the FCHL descriptor in a 100k $\Delta$-ML model gupta2021revving. Notably, with the aBoB-RBF(4) descriptor, the MAE already reaches 1.46 ppm using only 32k training entries, and approaches the 1 ppm accuracy threshold required for structure elucidation when trained on a 100k set. These results demonstrate the strong applicability of the continuous aBoB-RBF(4) descriptor for $\Delta$-ML modeling of NMR properties.
The aBoB-RBF($n$) descriptor was further evaluated for modeling 15N NMR chemical shifts for the nitrogen-containing molecules in the QM9NMR dataset. As $n$ increases from 0 to 5, the MAEs decrease to 4.08, 3.44, 2.99, 2.78, 2.73, and 2.75 ppm, respectively, using a training set of 100k molecules and an out-of-sample test set of 30k molecules. Using the same number of data points, Shiota et al. report an MAE of 3.42 ppm with the M3GNet GNN-TL embedding descriptor shiota2024universal. The systematic decrease in MAE with increasing $n$ indicates that the aBoB-RBF descriptors effectively capture the local atomic and electrostatic environments of nitrogen atoms in these molecules. Notably, the MAEs for 15N converge at $n=3$, whereas for 13C they converge at $n=4$, which can be rationalized in terms of differences in typical coordination numbers or valencies of the two atom types.
Overall, aBoB-RBF(4), with an MAE of 1.69 ppm, surpasses the previously reported benchmarks of 1.88 ppm for FCHL gupta2021revving and 1.87 ppm for the MACE-OFF23-small descriptor shiota2024universal.
Figure 6: Example of predicting 13C-NMR chemical shifts using mlqm9nmr module mlqm9nmr. For more detail, see mlqm9nmr module in SI.
Figure 7: DFT calculated (NMR: mPW1PW91/6-311+G(2d,p), geometry: B3LYP/6-31G(2df,p)) and ML-predicted 13C-NMR spectra for randomly chosen molecules, each from Drug12, Drug40, pyrimidinone, and GDBm validation set. ML predictions were made with mlqm9nmr mlqm9nmr using aBob-RBF(4) descriptor generated with PM7 geometries. See Figure 6 for an example query and SI for further details.